OpenStudy (anonymous):

The zeros of a polynomial are 3/4 and 1/2. Find the polynomial.

OpenStudy (anonymous):

Hint! you should find the following (x-3/4)(x-1/2)=0

OpenStudy (anonymous):

x^(2) − 10x + 3?

OpenStudy (anonymous):

i wouldn't use that although it would certainly work

OpenStudy (anonymous):

\((x-\frac{3}{4})\) gives a zero of \(\frac{3}{4}\) but so does \[(4x-3)\]

OpenStudy (anonymous):

would my answer be correct?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

you have to multiply out the factors that will give zeros at \(\frac{3}{4}\) and \(\frac{1}{2}\) as i wrote, \((4x-3)\) will give a zero at \(\frac{3}{4}\) because if you set \[4x-3=0\] you get \[4x=3\]\[x=\frac{3}{4}\]

OpenStudy (anonymous):

now you need the other factor, the one that will give a 0 at \(x=\frac{1}{2}\) so start by thinking "what would i set equal to zero and solve that would give an answer of \(\frac{1}{2}\)?

OpenStudy (anonymous):

sorry man im struggling on this one is this correct ? 4x^(2) − 10x + 3

OpenStudy (anonymous):

ok one factor is \(4x-3\) what is the other factor?

OpenStudy (anonymous):

i will tell you if it is that confusing

OpenStudy (anonymous):

(x-2) and ya i would like to know the answer.

OpenStudy (anonymous):

no actually \(x=2=0\iff x=2\) you want \(\frac{1}{2}\) so try \((2x-1)\)

OpenStudy (anonymous):

here is another way. x^2-Sx+p=0 where S is sum of the roots P =product of the roots S=3/4+1/2=5/4 P=(3/4)*(1/2)=3/8 so \[\Large x^2-\frac{5}{4}x+\frac{3}{8}=0\] simplify it.

OpenStudy (anonymous):

if you set \[2x-1=0\] you get \[2x=1\] \[x=\frac{1}{2}\] so the two factors are \((4x-3)\) and \((2x-1)\) and your last job is to multiply \[(4x-3)(2x-1)\]

OpenStudy (anonymous):

8x2 -10x +3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thats the one!

OpenStudy (anonymous):

whew hopefully now that you have seen it it will not be so hard in the future right?

OpenStudy (anonymous):

right thanks alot.

OpenStudy (anonymous):

if one zero is \(\frac{5}{7}\) we can say one factor is \((7x-5)\)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

hey can you help me with one more

OpenStudy (anonymous):

a different type of problem though

OpenStudy (anonymous):

sure go ahead ok then post in a new thread, better that way

OpenStudy (anonymous):

solve for w.